\begin{thebibliography}{63}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Abdou et~al.(2021)Abdou, Kulmizev, Hershcovich, Frank, Pavlick, and
  S{\o}gaard}]{abdou2021can}
Mostafa Abdou, Artur Kulmizev, Daniel Hershcovich, Stella Frank, Ellie Pavlick,
  and Anders S{\o}gaard. 2021.
\newblock Can language models encode perceptual structure without grounding?
  {A} case study in color.
\newblock \emph{arXiv preprint arXiv:2109.06129}.

\bibitem[{Altmann and Kamide(1999)}]{altmann1999incremental}
Gerry~TM Altmann and Yuki Kamide. 1999.
\newblock Incremental interpretation at verbs: Restricting the domain of
  subsequent reference.
\newblock \emph{Cognition}, 73(3):247--264.

\bibitem[{Barsalou(1983)}]{barsalou1983ad}
Lawrence~W Barsalou. 1983.
\newblock Ad hoc categories.
\newblock \emph{Memory \& cognition}, 11(3):211--227.

\bibitem[{Bender et~al.(2021)Bender, Gebru, McMillan-Major, and
  Shmitchell}]{bender2021dangers}
Emily~M Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret
  Shmitchell. 2021.
\newblock On the dangers of stochastic parrots: Can language models be too big?
\newblock In \emph{Proceedings of the 2021 ACM Conference on Fairness,
  Accountability, and Transparency}, pages 610--623.

\bibitem[{Bender and Koller(2020)}]{bender-koller-2020-climbing}
Emily~M. Bender and Alexander Koller. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.463} {Climbing
  towards {NLU}: {On} meaning, form, and understanding in the age of data}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 5185--5198, Online. Association for
  Computational Linguistics.

\bibitem[{Bisk et~al.(2020)Bisk, Holtzman, Thomason, Andreas, Bengio, Chai,
  Lapata, Lazaridou, May, Nisnevich et~al.}]{bisk2020experience}
Yonatan Bisk, Ari Holtzman, Jesse Thomason, Jacob Andreas, Yoshua Bengio, Joyce
  Chai, Mirella Lapata, Angeliki Lazaridou, Jonathan May, Aleksandr Nisnevich,
  et~al. 2020.
\newblock Experience grounds language.
\newblock \emph{arXiv preprint arXiv:2004.10151}.

\bibitem[{Block(1998)}]{block1998conceptual}
Ned Block. 1998.
\newblock Semantics, conceptual role.
\newblock \emph{Routledge encyclopedia of philosophy}, 8:652--657.

\bibitem[{Bonawitz et~al.(2012)Bonawitz, van Schijndel, Friel, and
  Schulz}]{bonawitz2012children}
Elizabeth~Baraff Bonawitz, Tessa~JP van Schijndel, Daniel Friel, and Laura
  Schulz. 2012.
\newblock Children balance theories and evidence in exploration, explanation,
  and learning.
\newblock \emph{Cognitive psychology}, 64(4):215--234.

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal,
  Neelakantan, Shyam, Sastry, Askell et~al.}]{brown2020language}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla
  Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell,
  et~al. 2020.
\newblock Language models are few-shot learners.
\newblock \emph{Advances in neural information processing systems},
  33:1877--1901.

\bibitem[{Casasanto and Lupyan(2015)}]{casasanto2015all}
Daniel Casasanto and Gary Lupyan. 2015.
\newblock {All Concepts Are Ad Hoc Concepts}.
\newblock In E.~Margolis and S.~Laurence, editors, \emph{The Conceptual Mind:
  New directions in the study of concepts}, pages 543--566. Cambridge: MIT
  Press.

\bibitem[{Dasgupta et~al.(2022)Dasgupta, Lampinen, Chan, Creswell, Kumaran,
  McClelland, and Hill}]{dasgupta2022language}
Ishita Dasgupta, Andrew~K Lampinen, Stephanie~CY Chan, Antonia Creswell,
  Dharshan Kumaran, James~L McClelland, and Felix Hill. 2022.
\newblock Language models show human-like content effects on reasoning.
\newblock \emph{arXiv preprint arXiv:2207.07051}.

\bibitem[{De~Deyne et~al.(2021)De~Deyne, Navarro, Collell, and
  Perfors}]{de2021visual}
Simon De~Deyne, Danielle~J Navarro, Guillem Collell, and Andrew Perfors. 2021.
\newblock Visual and affective multimodal models of word meaning in language
  and mind.
\newblock \emph{Cognitive Science}, 45(1):e12922.

\bibitem[{Deacon(1998)}]{deacon1998symbolic}
Terrence~William Deacon. 1998.
\newblock \emph{The symbolic species: The co-evolution of language and the
  brain}.
\newblock 202. WW Norton \& Company.

\bibitem[{Devlin et~al.(2018)Devlin, Chang, Lee, and
  Toutanova}]{devlin2018bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018.
\newblock Bert: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock \emph{arXiv preprint arXiv:1810.04805}.

\bibitem[{Elman(1990)}]{elman1990finding}
Jeffrey~L Elman. 1990.
\newblock Finding structure in time.
\newblock \emph{Cognitive science}, 14(2):179--211.

\bibitem[{Elman(2004)}]{elman2004alternative}
Jeffrey~L Elman. 2004.
\newblock An alternative view of the mental lexicon.
\newblock \emph{Trends in cognitive sciences}, 8(7):301--306.

\bibitem[{Fodor and Pylyshyn(1988)}]{fodor1988connectionism}
Jerry~A Fodor and Zenon~W Pylyshyn. 1988.
\newblock Connectionism and cognitive architecture: A critical analysis.
\newblock \emph{Cognition}, 28(1-2):3--71.

\bibitem[{Goldstein et~al.(2022)Goldstein, Zada, Buchnik, Schain, Price,
  Aubrey, Nastase, Feder, Emanuel, Cohen et~al.}]{goldstein2022shared}
Ariel Goldstein, Zaid Zada, Eliav Buchnik, Mariano Schain, Amy Price, Bobbi
  Aubrey, Samuel~A Nastase, Amir Feder, Dotan Emanuel, Alon Cohen, et~al. 2022.
\newblock Shared computational principles for language processing in humans and
  deep language models.
\newblock \emph{Nature neuroscience}, 25(3):369--380.

\bibitem[{Goodman et~al.(2011)Goodman, Ullman, and
  Tenenbaum}]{goodman2011learning}
Noah~D Goodman, Tomer~D Ullman, and Joshua~B Tenenbaum. 2011.
\newblock Learning a theory of causality.
\newblock \emph{Psychological review}, 118(1):110.

\bibitem[{Gopnik et~al.(1999)Gopnik, Meltzoff, and Kuhl}]{gopnik1999scientist}
Alison Gopnik, Andrew~N Meltzoff, and Patricia~K Kuhl. 1999.
\newblock \emph{The scientist in the crib: Minds, brains, and how children
  learn.}
\newblock William Morrow \& Co.

\bibitem[{Gopnik and Schulz(2004)}]{gopnik2004mechanisms}
Alison Gopnik and Laura Schulz. 2004.
\newblock Mechanisms of theory formation in young children.
\newblock \emph{Trends in cognitive sciences}, 8(8):371--377.

\bibitem[{Grand et~al.(2022)Grand, Blank, Pereira, and
  Fedorenko}]{grand2022semantic}
Gabriel Grand, Idan~Asher Blank, Francisco Pereira, and Evelina Fedorenko.
  2022.
\newblock Semantic projection recovers rich human knowledge of multiple object
  features from word embeddings.
\newblock \emph{Nature Human Behaviour}, pages 1--13.

\bibitem[{Hale(2001)}]{hale2001probabilistic}
John Hale. 2001.
\newblock A probabilistic earley parser as a psycholinguistic model.
\newblock In \emph{Second meeting of the north american chapter of the
  association for computational linguistics}.

\bibitem[{Hase et~al.(2021)Hase, Diab, Celikyilmaz, Li, Kozareva, Stoyanov,
  Bansal, and Iyer}]{hase2021language}
Peter Hase, Mona Diab, Asli Celikyilmaz, Xian Li, Zornitsa Kozareva, Veselin
  Stoyanov, Mohit Bansal, and Srinivasan Iyer. 2021.
\newblock Do language models have beliefs? methods for detecting, updating, and
  visualizing model beliefs.
\newblock \emph{arXiv preprint arXiv:2111.13654}.

\bibitem[{Hill et~al.(2017)Hill, Cho, Jean, and
  Bengio}]{hill2017representational}
Felix Hill, Kyunghyun Cho, S{\'e}bastien Jean, and Yoshua Bengio. 2017.
\newblock The representational geometry of word meanings acquired by neural
  machine translation models.
\newblock \emph{Machine Translation}, 31(1):3--18.

\bibitem[{Hill et~al.(2016)Hill, Cho, and Korhonen}]{hill2016learning}
Felix Hill, Kyunghyun Cho, and Anna Korhonen. 2016.
\newblock Learning distributed representations of sentences from unlabelled
  data.
\newblock \emph{arXiv preprint arXiv:1602.03483}.

\bibitem[{Jiang et~al.(2021)Jiang, Araki, Ding, and Neubig}]{jiang2021can}
Zhengbao Jiang, Jun Araki, Haibo Ding, and Graham Neubig. 2021.
\newblock How can we know when language models know? on the calibration of
  language models for question answering.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  9:962--977.

\bibitem[{Jumper et~al.(2021)Jumper, Evans, Pritzel, Green, Figurnov,
  Ronneberger, Tunyasuvunakool, Bates, {\v{Z}}{\'\i}dek, Potapenko
  et~al.}]{jumper2021highly}
John Jumper, Richard Evans, Alexander Pritzel, Tim Green, Michael Figurnov,
  Olaf Ronneberger, Kathryn Tunyasuvunakool, Russ Bates, Augustin
  {\v{Z}}{\'\i}dek, Anna Potapenko, et~al. 2021.
\newblock Highly accurate protein structure prediction with alphafold.
\newblock \emph{Nature}, 596(7873):583--589.

\bibitem[{Kadavath et~al.(2022)Kadavath, Conerly, Askell, Henighan, Drain,
  Perez, Schiefer, Dodds, DasSarma, Tran-Johnson, Johnston, El-Showk, Jones,
  Elhage, Hume, Chen, Bai, Bowman, Fort, Ganguli, Hernandez, Jacobson, Kernion,
  Kravec, Lovitt, Ndousse, Olsson, Ringer, Amodei, Brown, Clark, Joseph, Mann,
  McCandlish, Olah, and Kaplan}]{https://doi.org/10.48550/arxiv.2207.05221}
Saurav Kadavath, Tom Conerly, Amanda Askell, Tom Henighan, Dawn Drain, Ethan
  Perez, Nicholas Schiefer, Zac~Hatfield Dodds, Nova DasSarma, Eli
  Tran-Johnson, Scott Johnston, Sheer El-Showk, Andy Jones, Nelson Elhage,
  Tristan Hume, Anna Chen, Yuntao Bai, Sam Bowman, Stanislav Fort, Deep
  Ganguli, Danny Hernandez, Josh Jacobson, Jackson Kernion, Shauna Kravec,
  Liane Lovitt, Kamal Ndousse, Catherine Olsson, Sam Ringer, Dario Amodei, Tom
  Brown, Jack Clark, Nicholas Joseph, Ben Mann, Sam McCandlish, Chris Olah, and
  Jared Kaplan. 2022.
\newblock \href {https://doi.org/10.48550/ARXIV.2207.05221} {Language models
  (mostly) know what they know}.

\bibitem[{Kocijan et~al.(2020)Kocijan, Lukasiewicz, Davis, Marcus, and
  Morgenstern}]{kocijan2020review}
Vid Kocijan, Thomas Lukasiewicz, Ernest Davis, Gary Marcus, and Leora
  Morgenstern. 2020.
\newblock A review of winograd schema challenge datasets and approaches.
\newblock \emph{arXiv preprint arXiv:2004.13831}.

\bibitem[{Kriegeskorte et~al.(2008)Kriegeskorte, Mur, and
  Bandettini}]{kriegeskorte2008representational}
Nikolaus Kriegeskorte, Marieke Mur, and Peter~A Bandettini. 2008.
\newblock Representational similarity analysis-connecting the branches of
  systems neuroscience.
\newblock \emph{Frontiers in systems neuroscience}, page~4.

\bibitem[{Kumar et~al.(2022)Kumar, Sumers, Yamakoshi, Goldstein, Hasson,
  Norman, Griffiths, Hawkins, and Nastase}]{kumar2022reconstructing}
Sreejan Kumar, Theodore~R Sumers, Takateru Yamakoshi, Ariel Goldstein, Uri
  Hasson, Kenneth~A Norman, Thomas~L Griffiths, Robert~D Hawkins, and Samuel~A
  Nastase. 2022.
\newblock Reconstructing the cascade of language processing in the brain using
  the internal computations of a transformer-based language model.
\newblock \emph{bioRxiv}.

\bibitem[{Lake and Murphy(2021)}]{lake2021word}
Brenden~M Lake and Gregory~L Murphy. 2021.
\newblock Word meaning in minds and machines.
\newblock \emph{Psychological review}.

\bibitem[{Lample et~al.(2017)Lample, Conneau, Denoyer, and
  Ranzato}]{lample2017unsupervised}
Guillaume Lample, Alexis Conneau, Ludovic Denoyer, and Marc'Aurelio Ranzato.
  2017.
\newblock Unsupervised machine translation using monolingual corpora only.
\newblock \emph{arXiv preprint arXiv:1711.00043}.

\bibitem[{Levy(2008)}]{levy2008expectation}
Roger Levy. 2008.
\newblock Expectation-based syntactic comprehension.
\newblock \emph{Cognition}, 106(3):1126--1177.

\bibitem[{Lewkowycz et~al.(2022)Lewkowycz, Andreassen, Dohan, Dyer,
  Michalewski, Ramasesh, Slone, Anil, Schlag, Gutman-Solo
  et~al.}]{lewkowycz2022solving}
Aitor Lewkowycz, Anders Andreassen, David Dohan, Ethan Dyer, Henryk
  Michalewski, Vinay Ramasesh, Ambrose Slone, Cem Anil, Imanol Schlag, Theo
  Gutman-Solo, et~al. 2022.
\newblock Solving quantitative reasoning problems with language models.
\newblock \emph{arXiv preprint arXiv:2206.14858}.

\bibitem[{Li et~al.(2021)Li, Nye, and Andreas}]{li2021implicit}
Belinda~Z Li, Maxwell Nye, and Jacob Andreas. 2021.
\newblock Implicit representations of meaning in neural language models.
\newblock \emph{arXiv preprint arXiv:2106.00737}.

\bibitem[{Margolis and Laurence(1999)}]{margolis1999concepts}
Eric~Ed Margolis and Stephen~Ed Laurence. 1999.
\newblock \emph{Concepts: Core Readings.}
\newblock The MIT Press.

\bibitem[{McClelland et~al.(2020)McClelland, Hill, Rudolph, Baldridge, and
  Sch{\"u}tze}]{mcclelland2020placing}
James~L McClelland, Felix Hill, Maja Rudolph, Jason Baldridge, and Hinrich
  Sch{\"u}tze. 2020.
\newblock Placing language in an integrated understanding system: Next steps
  toward human-level performance in neural language models.
\newblock \emph{Proceedings of the National Academy of Sciences},
  117(42):25966--25974.

\bibitem[{Mikolov et~al.(2013)Mikolov, Yih, and Zweig}]{mikolov2013linguistic}
Tom{\'a}{\v{s}} Mikolov, Wen-tau Yih, and Geoffrey Zweig. 2013.
\newblock Linguistic regularities in continuous space word representations.
\newblock In \emph{Proceedings of the 2013 conference of the north american
  chapter of the association for computational linguistics: Human language
  technologies}, pages 746--751.

\bibitem[{Murphy and Medin(1985)}]{murphy1985role}
Gregory~L Murphy and Douglas~L Medin. 1985.
\newblock The role of theories in conceptual coherence.
\newblock \emph{Psychological review}, 92(3):289.

\bibitem[{Ortega et~al.(2021)Ortega, Kunesch, Del{\'e}tang, Genewein,
  Grau-Moya, Veness, Buchli, Degrave, Piot, Perolat et~al.}]{ortega2021shaking}
Pedro~A Ortega, Markus Kunesch, Gr{\'e}goire Del{\'e}tang, Tim Genewein, Jordi
  Grau-Moya, Joel Veness, Jonas Buchli, Jonas Degrave, Bilal Piot, Julien
  Perolat, et~al. 2021.
\newblock Shaking the foundations: delusions in sequence models for interaction
  and control.
\newblock \emph{arXiv preprint arXiv:2110.10819}.

\bibitem[{Packard et~al.(1980)Packard, Crutchfield, Farmer, and
  Shaw}]{packard1980geometry}
Norman~H Packard, James~P Crutchfield, J~Doyne Farmer, and Robert~S Shaw. 1980.
\newblock Geometry from a time series.
\newblock \emph{Physical review letters}, 45(9):712.

\bibitem[{Patel and Pavlick(2021)}]{patel2021mapping}
Roma Patel and Ellie Pavlick. 2021.
\newblock Mapping language models to grounded conceptual spaces.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Pavlick(2022)}]{pavlick2022semantic}
Ellie Pavlick. 2022.
\newblock Semantic structure in deep learning.
\newblock \emph{Annual Review of Linguistics}, 8:447--471.

\bibitem[{Peters et~al.(2018)Peters, Neumann, Iyyer, Gardner, Clark, Lee, and
  Zettlemoyer}]{peters2018deep}
ME~Peters, M~Neumann, M~Iyyer, M~Gardner, C~Clark, K~Lee, and L~Zettlemoyer.
  2018.
\newblock Deep contextualized word representations. arxiv 2018.
\newblock \emph{arXiv preprint arXiv:1802.05365}, 12.

\bibitem[{Piantadosi(2021)}]{piantadosi2021computational}
Steven~T Piantadosi. 2021.
\newblock The computational origin of representation.
\newblock \emph{Minds and machines}, 31(1):1--58.

\bibitem[{Pierce(2002)}]{pierce2002types}
Benjamin~C Pierce. 2002.
\newblock \emph{Types and programming languages}.
\newblock MIT press.

\bibitem[{Putnam(1974)}]{putnam1974meaning}
Hilary Putnam. 1974.
\newblock Meaning and reference.
\newblock \emph{The journal of philosophy}, 70(19):699--711.

\bibitem[{Quine(1977)}]{quine1977natural}
W.V.O. Quine. 1977.
\newblock Natural kinds.
\newblock In S.P. Schwartz, editor, \emph{Naming, necessity, and natural
  kinds}, pages 155--175. Ithaca, NY: Cornell University Press.

\bibitem[{Rule et~al.(2020)Rule, Tenenbaum, and Piantadosi}]{rule2020child}
Joshua~S Rule, Joshua~B Tenenbaum, and Steven~T Piantadosi. 2020.
\newblock The child as hacker.
\newblock \emph{Trends in cognitive sciences}, 24(11):900--915.

\bibitem[{Russell(2010)}]{russell2010artificial}
Stuart~J Russell. 2010.
\newblock \emph{Artificial intelligence a modern approach}.
\newblock Pearson Education, Inc.

\bibitem[{Santoro et~al.(2021)Santoro, Lampinen, Mathewson, Lillicrap, and
  Raposo}]{santoro2021symbolic}
Adam Santoro, Andrew Lampinen, Kory Mathewson, Timothy Lillicrap, and David
  Raposo. 2021.
\newblock Symbolic behaviour in artificial intelligence.
\newblock \emph{arXiv preprint arXiv:2102.03406}.

\bibitem[{Schrimpf et~al.(2021)Schrimpf, Blank, Tuckute, Kauf, Hosseini,
  Kanwisher, Tenenbaum, and Fedorenko}]{schrimpf2021neural}
Martin Schrimpf, Idan~Asher Blank, Greta Tuckute, Carina Kauf, Eghbal~A
  Hosseini, Nancy Kanwisher, Joshua~B Tenenbaum, and Evelina Fedorenko. 2021.
\newblock The neural architecture of language: Integrative modeling converges
  on predictive processing.
\newblock \emph{Proceedings of the National Academy of Sciences},
  118(45):e2105646118.

\bibitem[{Shepard(1980)}]{shepard1980multidimensional}
Roger~N Shepard. 1980.
\newblock Multidimensional scaling, tree-fitting, and clustering.
\newblock \emph{Science}, 210(4468):390--398.

\bibitem[{Shepard and Chipman(1970)}]{shepard1970second}
Roger~N Shepard and Susan Chipman. 1970.
\newblock Second-order isomorphism of internal representations: Shapes of
  states.
\newblock \emph{Cognitive psychology}, 1(1):1--17.

\bibitem[{Takens(1981)}]{takens1981detecting}
Floris Takens. 1981.
\newblock Detecting strange attractors in turbulence.
\newblock In \emph{Dynamical systems and turbulence, Warwick 1980}, pages
  366--381. Springer.

\bibitem[{Tenenbaum et~al.(2011)Tenenbaum, Kemp, Griffiths, and
  Goodman}]{tenenbaum2011grow}
Joshua~B Tenenbaum, Charles Kemp, Thomas~L Griffiths, and Noah~D Goodman. 2011.
\newblock How to grow a mind: Statistics, structure, and abstraction.
\newblock \emph{science}, 331(6022):1279--1285.

\bibitem[{Ullman et~al.(2012)Ullman, Goodman, and Tenenbaum}]{ullman2012theory}
Tomer~D Ullman, Noah~D Goodman, and Joshua~B Tenenbaum. 2012.
\newblock Theory learning as stochastic search in the language of thought.
\newblock \emph{Cognitive Development}, 27(4):455--480.

\bibitem[{Vuli{\'c} et~al.(2017)Vuli{\'c}, Gerz, Kiela, Hill, and
  Korhonen}]{vulic2017hyperlex}
Ivan Vuli{\'c}, Daniela Gerz, Douwe Kiela, Felix Hill, and Anna Korhonen. 2017.
\newblock Hyperlex: A large-scale evaluation of graded lexical entailment.
\newblock \emph{Computational Linguistics}, 43(4):781--835.

\bibitem[{Wason and Johnson-Laird(1972)}]{wason1972psychology}
Peter~Cathcart Wason and Philip~Nicholas Johnson-Laird. 1972.
\newblock \emph{Psychology of reasoning: Structure and content}, volume~86.
\newblock Harvard University Press.

\bibitem[{Wittgenstein(1953)}]{wittgenstein2010philosophical}
Ludwig Wittgenstein. 1953.
\newblock \emph{Philosophical investigations}.
\newblock John Wiley \& Sons.

\bibitem[{Xu et~al.(2020)Xu, Patwary, Shoeybi, Puri, Fung, Anandkumar, and
  Catanzaro}]{xu2020megatron}
Peng Xu, Mostofa Patwary, Mohammad Shoeybi, Raul Puri, Pascale Fung, Anima
  Anandkumar, and Bryan Catanzaro. 2020.
\newblock Megatron-cntrl: Controllable story generation with external knowledge
  using large-scale language models.
\newblock \emph{arXiv preprint arXiv:2010.00840}.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[]{acl}
\usepackage{times}
\usepackage{latexsym}

\usepackage{type1cm}
\usepackage{lettrine}
\usepackage{url}
\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\title{Meaning without reference in large language models}

\author{Steven T. Piantadosi \\
       Department of Psychology \\
       Helen Wills Neuroscience Institute\\
       University of California, Berkeley \\   stp@berkeley.edu
       \And
        Felix Hill \\ 
       DeepMind \\
        felixhill@deepmind.com }

\begin{document}

\maketitle

\begin{abstract}
The widespread success of large language models (LLMs) has been met with skepticism that they possess anything like human concepts or meanings. Contrary to claims that LLMs possess no meaning whatsoever, we argue that they likely capture important aspects of meaning, and moreover work in a way that approximates a compelling account of human cognition in which meaning arises from \emph{conceptual role}. Because conceptual role is defined by the relationships between internal representational states, meaning cannot be determined from a model's architecture, training data, or objective function, but only by examination of how its internal states relate to each other. This approach may clarify why and how LLMs are so successful and suggest how they can be made more human-like. 
\end{abstract}

\vspace{15pt}
\noindent \lettrine{L}{arge} language models (LLMs) have begun to display an array of competencies that were long thought to be out of the reach of neural networks. At the same time, critics have been vocal that existing methods, model structures, and training paradigms will be insufficient for anything like human language use. In particular, it is argued that LLMs will never achieve “meaning” or “understanding” due to either the objective function they optimize, the format of their internal representations, or the type of training data that they receive. This short comment aims to contest the claims that LLMs are necessarily incapable of acquiring meaning, and suggest that LLM meanings may already be similar to human-like meaning in many (but not all) ways. We argue that LLMs have likely already achieved some key aspects of meaning which, while imperfect, mirror the way meanings work in cognitive theories, as well as approaches to the philosophy of language. 

LLMs are trained to predict words from massive datasets of text from the internet. They typically contain billions of parameters that are jointly optimized---at great computational, energy, and financial expense \citep{bender2021dangers}---to make predictions about word occurrence from surrounding context. This setup differs from human language acquisition in data scale and format. On the other hand, the core objective of word prediction is a central piece of human language processing \citep[e.g.][]{altmann1999incremental,hale2001probabilistic,levy2008expectation} and has long been shown capable of providing a learning signal from which linguistic structures and semantic categories can emerge \citep{elman1990finding}.

A key debate in recent literature has been about \emph{what} models trained in this manner come to know. A prominent view is that models trained only on text cannot acquire realistic meanings because they lack reference, or connection to objects in the real world. \citet{bender-koller-2020-climbing} illustrate this with an ``octopus test.'' They imagine an octopus that learns to use words correctly by eavesdropping on a conversation between two people on land. If the octopus has no access to the referents of the words then there are gaps in its meaning for the words. For example, if the octopus must suddenly determine which \emph{object} is a coconut, then its expertise in using the word ``coconut'' won’t help. Its knowledge of co-occurrence statistics between ``coconut'' and other words won't help either since it also knows nothing about the referents of other words. The octopus simply does not have the required meanings to find the coconut. Humans learners don't have this problem because their input---and consequently representations---are tied to real-world referents. Bender \& Koller's position is that no amount of predictive linguistic savvy can give models that are trained on text alone the knowledge of reference they need to acquire meanings. 

\paragraph{Meaning and reference}

The octopus test assumes that reference determines meaning, but in fact cognitive scientists and philosophers have found a variety of problems with this view. One is that there are many terms that are meaningful to us but have no discernible referent at all, such as abstract words like “justice” and ``wit.'' People can think of new concepts like “aphid-sized accordion” that don’t exist (thus no referent), or even terms that have no \emph{possible} referent like “perpetual motion machine” or ``imaginary cup of tea.'' We can think of concepts like “king of San Francisco” that pick out nobody, but are at least meaningful enough to reason about (for example: “If there was a King of San Francisco, he’d live in The Presidio.”). Other examples show reference can be quite decoupled from meaning. Terms like “treaty” and “contract” are often thought to have a concrete referent, but what is important is actually an abstract entity: a treaty is still valid if the piece of paper is destroyed. Frege’s example is of the “morning star” and the “evening star.” Both are terms for the planet Venus, but were once conceived of as  different entities without knowing that they are the same object. 

This problem is not solely an issue with abstract concepts. Many concrete terms seem not to get their most important semantic properties from reference either. Consider an example like a “postage stamp.” Everyone can conjure up an image of a typical, physical, postage stamp. But with further scrutiny, none of the concrete features of typical stamps seem necessary. We can imagine a country whose postage stamps were made of clear glass, for example, or stamps that were microscopic so that they could not be seen, or stamps that were paid for and tracked entirely online, or others that were larger than a house (for mailing very large packages). We could think of postage stamps that were \emph{RFID} tags that went inside the envelope. Maybe intelligent ants would use postage stamps that were pheromones rather than paper; in the future we might have postage stamps that sprout wings and fly your letter to its intended location. Already we have stamps that only have barcodes and stamps that you can draw yourself. You know the term “postage stamp” but there is no way that you could have considered all of the possible referents yet, so reference cannot be what determines the concept. 

Similar arguments were made by \citet{wittgenstein2010philosophical} when considering the concept of ``water'', leading him to conclude that reference plays little or no role in determining meaning in the general case. One way to explain these observations is to assume that our meaning for terms like ``postage stamp'' (or ``water'') may be primarily determined by the role these concepts play in some greater mental theory. Very roughly, people call something a ``postage stamp'' if you pay for it and then attach it to a letter in order to have the letter to be delivered. In this view, the meaning of the word is intrinsically intertwined with other concepts like “payment”, “letter” and “delivery.” The interrelation is key \cite[see, e.g.][]{deacon1998symbolic,santoro2021symbolic} because when the associated terms shift in meaning even slightly (e.g. credit cards were developed as wholly new a form of “payment”) the meaning of “postage stamp” comes along for the ride (we know right away that they can be paid for by a credit card). Such relationships between concepts are \emph{the} essential, defining, aspects of meaning and, in fact, possessing the appropriate relationships allows you to determine the reference. This view makes sense of the puzzling examples above.

\paragraph{Conceptual role theory}

In philosophy of mind, versions of this approach go by the name ``conceptual role theory.'' Following \citet{block1998conceptual}, consider statements in physics like $f=m\cdot a$. This equation is not exactly a definition of $f$ (\emph{force}), nor is it a definition of \emph{mass} ($m$) or \emph{acceleration} ($a$), but, is a statement about the interrelationship of $f$, $m$, and $a$. Most of us can't give much more detail about the ultimate physical reference of these terms---forces have something to do with interacting elementary particles (whatever those are), masses have something to do with the Higgs field (whatever that is)---but our thoughts about $f=m\cdot a$ certainly do not seem meaningless. Many believe that conceptual roles are one of the most promising ways to characterize human concepts \emph{in general} \citep[for an overview of this and competing theories, see, e.g., ][]{margolis1999concepts}. 
\citet{murphy1985role} for example argue that our organization of categories is based on entire theories of structured conceptual domains (rather than simple features or similarities), an idea they trace to \citet{quine1977natural}. Murphy and Medin note how we might reflexively consider a composite category such as “prime numbers or apples” to be an unnatural or incoherent set of entities. However, if we know someone called Wilma who is a number theorist grew up on an apple farm, then the category “topics of conversation with Wilma”, involving the same constituent entities, seems perfectly reasonable. What is a natural category or concept depends on our mental conception of how the underlying pieces relate, and concepts can even be assembled fluidly, in an ad hoc manner or context-dependent manner \citep{barsalou1983ad,casasanto2015all}.
Theories in cognition have also been explored in learning models \citep[e.g.][]{goodman2011learning,ullman2012theory} and experimentally shown to shape how children explore the world \citep[e.g.][]{gopnik1999scientist,gopnik2004mechanisms,bonawitz2012children}. 

\paragraph{Conceptual role in LLMs}

If anything like this view is correct, then the search for meaning in learning models---or brains---should focus on understanding the way that the systems' internal representational states \emph{relate to each other}. Once a learning model finds it probable that “postage stamps” are “affixed” to “letters” so they can be “delivered,” then it has acquired some important pieces of conceptual role for these terms. It would not be possible to conclude anything about what meanings a system does and does not possess from its training data or architecture because these may not be informative about how the internal states relate to each other. 

Relations between internal states have been long emphasized in cognitive theories \citep{shepard1970second,deacon1998symbolic,fodor1988connectionism}, for example early attempts to discover the geometry of psychological space \citep{shepard1980multidimensional} and more recent analyses of brain data based on representational similarity \citep{kriegeskorte2008representational}. \citet{elman2004alternative} argues for a closely related view of the mental lexicon in which a word's meaning is the effect it has on other mental states. In deep learning models, the relational geometry of vector representations have been examined for instance in analogy problems~\citep{mikolov2013linguistic}, match to human similarities~\citep{hill2017representational}, and encoding of humanlike gradient distinctions~\citep{vulic2017hyperlex}. \citet{grand2022semantic} show that semantic embeddings from these models capture gradient scales of multiple features, like from ``small'' to ``big'' or ``safe'' to ``dangerous.'' It is even possible to align the word representations acquired by text-based models across languages to translate between them effectively with no prior knowledge of which words or phrases should have the same meaning~\citep{lample2017unsupervised}.  Similarly, \citet{abdou2021can} show that a model trained on text can recover key geometry of color space, even without grounding; with a few examples of grounding, LLMs are able to align their structure with the real grounded one, suggesting that they already possess the right relations \cite{patel2021mapping}. Importantly, as the performance of LLMs has improved in recent years the extent to which their relational geometry reflects human data has also consistently increased~\citep{peters2018deep,devlin2018bert,brown2020language}. Larger models also better reflect the human tendency for semantic or mental models to influence formal reasoning behaviour~\citep{wason1972psychology}, on challenging logic problems that are not observed in their training data~\citep{dasgupta2022language}. Moreover, this increasing correspondence between LLMs and human data is not observed only behaviourally. Recent fMRI studies show that the semantic models that best account for the representational geometry and processing activity of human brains are precisely the neural network LLMs which are trained on the largest amount of data \citep{schrimpf2021neural,goldstein2022shared,kumar2022reconstructing}. 

Many of the tasks that LLMs succeed on are ones that require maintaining the right relationships between concepts. Impressively, the largest models can now devise coherent narratives~\citep{brown2020language}, extend stories~\citep{xu2020megatron,li2021implicit}, answer factual questions~\citep{jiang2021can} solve Winograd Schema~\citep{kocijan2020review} and resolve complex quantitative reasoning problems~\citep{lewkowycz2022solving}. Increasingly, such models are even aware of the likelihood that they can answer a given question correctly; i.e. they have an explicit sense of the extent of their own knowledge~\citep{https://doi.org/10.48550/arxiv.2207.05221}. Each of these capacities requires, in some way or another, sensitivity to conceptual roles because the required words and concepts must be used jointly together in a coherent way that mimics how humans would. 

Despite these empirical successes, there are many places where these models can still be improved (for detailed analyses, see, \citet{lake2021word,mcclelland2020placing,pavlick2022semantic}). \citet{lake2021word} emphasize the need for reference, inference, better and more robust compositionality, more structure and more consistent abstract reasoning. Models trained on multimodal datasets show better match to human judgements than those trained on text alone \citep{hill2016learning,de2021visual}; at the same time, even multimodal LLMs are missing many aspects of a complete theory of semantics, including the ability to simulate situations in which their physical or linguistic behaviour affects their environment \citep{mcclelland2020placing} as well as knowledge of the goals and desires that drive how people use words \citep{bisk2020experience,lake2021word}.

Our claim, then, is not that LLMs perfectly capture human concepts or perfectly reflect human meaning. Unlike the more radical perspectives entertained by Wittgenstein or Quine, we also do not consider that reference should play \emph{no} role in a principled treatment of meaning. Instead, we find it productive to consider reference as just one (optional) aspect of a word's full conceptual role. It is relevant for some concepts \citep[see][]{putnam1974meaning} and not others---just like color, valence, or teleology is relevant for some concepts and not others. Experience of both agency and a perceptual environment similar to our own may lead to the richest, most human-like understanding of language in machines \citep{bisk2020experience,mcclelland2020placing}. 

As these improvements are made, the models will come into closer alignment with humans, and each such improvement will enrich the model's sense of meaning. This process of progressive enrichment is also found in human concepts. When people discovered that $H_2O$ was the chemical composition of water, they grew their conceptual network and even revised their reference for the term. There was no hard transition from a meaningless concept of ``water'' to a  meaningful one. Some meaning was there all along because ``water'' had a conceptual role even before its chemical composition was known. What changed was the richness and interconnection of this concept---the way in which it was related to other concepts like ``hydrogen'' and ``oxygen.'' In much the same way, we see no reason to assume that the world of a system that receives input from a single sensory modality is meaningless, even if the addition of further sensors provides clear enrichment. When thinking about improving LLMs it we should therefore consider ways to enrich the internal conceptual roles of these systems, including to better reflect the structure, inference and algorithmic sophistication of humans \citep{tenenbaum2011grow,lake2021word,rule2020child}. 

\paragraph{Discovering conceptual role}

Conceptual role theory also provides a compelling way to understand learning, including the way in which neural networks may come to represent symbolic processes. A symbol like $AND$ only means logical conjunction if it interacts (composes) with others symbols like $TRUE$ and $FALSE$ in the appropriate way---i.e. when it has the right conceptual role in the broader system of symbols. The technique of \emph{church-encoding} in mathematical logic \cite[see][]{pierce2002types} provides a way to understand how such roles may be learned within neural networks or dynamical systems \cite{piantadosi2021computational}. In church-encoding, a representation is constructed in one system (e.g. lambda calculus or a neural network) in order to mimic the behavior of another system (e.g. boolean logic) in the sense that the representations in the first system interact with each other in a way that yields the desired conceptual roles of the second. \citet{piantadosi2021computational} shows how a church-encoding learner could acquire structures like logic, lists, trees, hierarchies, numbers, quantifiers, and recursion without possessing them to start, and how this metaphor may provide an ``assembly language'' that translates from symbolic computational or cognitive theories into underlying implementations. In this view, neural networks would train their parameters so that their internal, intrinsic dynamics church-encode the conceptual roles of a targeted domain.

The key question for LLMs is whether training to predict text could actually support discovery of conceptual roles. To us the question is empirical, and we believe has been answered in a promising, partial affirmative by studies showing success on tasks that require knowledge of relationships between concepts. Text provides such clues to conceptual role because human conceptual roles generated the text. Analogously, it is possible  to build a theory of gravity from measurements of the moon's movement because gravity \emph{generated} these movements; the goal of essentially all inductive learning techniques is to invert from observations to likely generating processes or parameters. Moreover, the entailment relationships between sentences are often intrinsically related to analogous patterns in thought \citep[e.g.][]{fodor1988connectionism}, meaning that a model which captures how sentences relate to each other might indeed capture how thoughts relate to each other. One helpful analogy is that of \emph{embedding theorems} in dynamical systems \cite[e.g.][]{packard1980geometry,takens1981detecting} which allow some properties of systems to be recovered from seemingly impoverished representations of their state. In the paper ``Geometry from a Time Series'', \citet{packard1980geometry} show, remarkably, that one can sometimes reconstruct the geometry of a multi-dimensional dynamical system from a \emph{one}-dimensional projection of its state. Thus, information about high-dimensional state (sometimes essentially all of it) can be decoded from the trajectory of low-dimensional projection. People use concepts in thinking and reasoning based on their meaning, and text is a low-dimensional projection of some of these patterns of use, so it is plausible that some properties of the real meaning could be inferred from text. At the very least, embedding theorems illustrate that there may not be a simple way to intuit what a learner can or cannot deduce about the underlying mental states from text alone.

The protein folding neural network AlphaFold~\citep{jumper2021highly} provides further evidence that transformer-based networks can infer and generalise complex latent multi-entity structures. AlphaFold is trained to predict the configuration of single proteins only, but acquires actionable knowledge of concepts not explicitly present in its training data. Despite never seeing a zinc ion, AlphaFold often perfectly infers its location and places all the protein side chains correctly right around it. Some proteins only fold with multiple copies of themselves (homomers). Again, AlphaFold has never seen more than one copy of a protein, but often infers both the number of required copies and their relative placement correctly. This suggests that the process of predicting the structure of single proteins enabled the AlphaFold network to infer non-trivial facts about chemistry and biology.
\paragraph{Communicative intentions}

Separable from meaning and reference, many have also rejected the idea that LLMs produce language with \emph{intention} \citep{bender2021dangers,bender-koller-2020-climbing}. Because LLMs are trained only on sequence prediction, they are argued to be, ``stochastic parrots'' \cite{bender2021dangers} or just sophisticated ``auto-complete'' algorithms\footnote{\tiny\url{https://garymarcus.substack.com/p/nonsense-on-stilts}} that cannot access the intentions of those who produced their training data, and themselves produce language without intending anything in particular. But in our view, a key difference between autocompleting parrots and LLMs is that the latter have rich, causal, and structured internal states. 

One view of intent is semantic, corresponding to whether the language they produce arises from an internal representation of (intended) meaning. The conversion of internal states into language and back \emph{is} the essential function of LLMs, embedded in their architecture and training. We have argued that LLM's internal state has some notions of conceptual role, so LLM's utterances have the semantic intent corresponding to these roles. 

Another view of intent is pragmatic, asking what might be achieved by producing a sentence. This corresponds to asking whether they engage in any goal-directed \emph{planning}~\citep{russell2010artificial} when producing language. We consider it probable that multi-layer LLMs do, in an emergent, implicit sense, execute a form of planning as part of the process of repeated (self-attention-based) analyses of current and past inputs. These computations likely involve representation of the current situation and at least implicit evaluation of consequences of utterances. Recent work has argued that LLMs posses model-like belief structures \cite{hase2021language} and update representations of dynamic semantics, objects and situations, throughout a discourse \cite{li2021implicit}. These emergent semantics causally determine LLM output. Of course, differences between humans and LLMs in their training experience and objectives mean that the planning process in LLMs is less explicit and sophisticated than in humans (making errors more likely, for instance, in cases of hypothetical or counterfactual reasoning~\citep{ortega2021shaking}). 
 
\paragraph{Conclusion}

Bender \& Koller argue that text-based LLMs will never have meaning because these models lack reference. However, they do not demonstrate that reference is the key to meaning---instead they assume it. As we have argued, this assumption is hard to reconcile with theories of cognition and the phenomena that motivate them. People are happy to think about concepts without referents and otherwise often don't know many details of reference. Meaning instead seems to come from the way concepts relate to \emph{each other}. It is these interrelations that LLMs know something about since their internal geometries and trajectories approximate those of humans. Like people who don't know that water is $H_2O$ and so could not pick it out based on chemical composition, Bender \& Koller's octopus lacks some aspects of conceptual role like physical appearance. But, both the octopus and people know other parts of conceptual role that are sophisticated in their own right. If theories about conceptual role are the correct account, then LLMs likely already share the foundation of how our own concepts get their meaning. 

\section*{Acknowledgements}
We are grateful to Ev Fedorenko, Adam Santoro, MH Tessler, Dileep George, John Hale, Miguel Figurnov, John Jumper, Dharsh Kumaran, Matt Botvinick, Paul Smolensky and 
Chris Dyer for detailed comments. 

\bibliography{acl_latex}




\end{document}